OpenStudy (anonymous):

answer the stand form of (3x-5) (3x+5) = 10 (x-3)quantity squared + 7 =0 2(x-1)quantity squared -6 =0 x(2x+7=(x+3) (x-3) -3x(x+5) = 8 (2x+3) quantity squared (x+3) quantitysquared .. JESA :)

OpenStudy (anonymous):

Is it 10(x-3) ^2..? Or the whole equation squared?

OpenStudy (anonymous):

Please writr properly using equation.

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

Is it (3x-5) (3x+5) = 10(x-3)^2+ 7= 0 ?

OpenStudy (anonymous):

yes .. wat is the answer ?

OpenStudy (anonymous):

Solve (3x-5)(3x+5)=0

OpenStudy (anonymous):

9x_25=0 .. ?

mathslover (mathslover):

yes jesa it is : \((3x-5)(3x+5)=0\) ===>\(9x^2-25\)

OpenStudy (anonymous):

how the others .. ?

mathslover (mathslover):

try to do them on ur own .. tell me where r u having problem

OpenStudy (anonymous):

plzzzzzzzz .

OpenStudy (anonymous):

uhhhhmm .. i dont know how to solve it !

OpenStudy (anonymous):

im sorry !

mathslover (mathslover):

can u write the second one in expression form ?

mathslover (mathslover):

i mean : (x-3)quantity squared + 7 =0 this into an expression

mathslover (mathslover):

did u get what i mean ?

OpenStudy (anonymous):

(x-3)2 + 7+=0

mathslover (mathslover):

ok very good ; \((x-3)^2+7=0\) Do you know how to solve \((x-3)^2\) ?

OpenStudy (anonymous):

x squared-9 /

mathslover (mathslover):

no .. @jesa just try to find it out with this identity : \(\huge{(a-b)^2=a^2+b^2-2ab}\)

OpenStudy (anonymous):

i dont know !

mathslover (mathslover):

put a = x , b = 3 can u do it now ?

OpenStudy (anonymous):

Okay Jesa.. \[(a-b)^{2}= (a-b)(a-b)\].. if you expnad you'll get \[a.a - a.b-b.a + b.b\], Which is equal to \[ a^{2}-2ab- b^{2}\]. Now can you solve?